• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 23:57:33 on July20,2006

    Entry number 174663

    keyword=Shift summary

    16:00 Target cool-down underway; going (and will continue to go) very slowly because of "issues" with insufficient cooling power from ESR.
    18:30 Hall locked up to let MCC begin trying to get beam in.
    19:45 MCC calls to say they will now try to bring in tuned beam.
    20:10 MCC is unable to read BPMs in Hall; try rebooting IOC, no luck; calling in expert.
    21:00 MCC got this working - then found that there was residual field in the Compton magnets...have to cycle those magnets. This will take time.
    21:15 Low-pressure alarm going off on target loop (loop 1, LH2) - this was an expected eventuality that needs a quick access to let in more hydrogen from the tank; also, the PID for the target heater isn't working. Meekins coming in to address both issues.
    22:00-23:00 Access to look at target issues. More hydrogren let in, low pressure alarm now fixed. However, the problem with heater PID remains unfixed (see halog entry on this by Craver).
    23:10 Back to beam permit. MCC working on ensuring correct optics now into the hall.